If Helga Inc. issued a bond that is currently selling for $950 has 7 years left until maturity and currently as a 10.4% yield to maturity. What must the bond’s coupon price be?

$109.21

$93.59

$301.83

$46.88

None of these
Answer #1

Current value=Annual coupon*Present value of annuity factor(10.4%,7)+$1000*Present value of discounting factor(10.4%,7)

950=Annual coupon*4.80496295+1000*0.500283853

Annual coupon=(950-500.283853)/4.80496295

=$93.59(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=Annual coupon[1-(1.104)^-7]/0.104

=Annual coupon*4.80496295

2.Present value of discounting factor=1000/1.104^7

=1000*0.500283853